# references.bib

### v3

@@ Abstract
With the advent of artificial intelligence (AI) and machine learning (ML), various domains of science and engineering communities have leveraged data-driven surrogates to model complex systems from numerous sources of information (data). The proliferation has led to significant reduction in cost and time involved in development of superior systems designed to perform specific functionalities. A high proposition of such surrogates is built extensively fusing multiple sources of data, may it be published papers, patents, open repositories, or other resources. However, not much attention has been paid to the differences in quality and comprehensiveness of the known and unknown underlying physical parameters of the information sources that could have downstream implications during system optimization. Towards resolving this issue, a multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed. The individual data sources are tagged as a characteristic categorical variable that are mapped into a physically interpretable latent space, allowing the development of source-aware data fusion modeling. Additionally, a dissimilarity metric based on the latent variables of LVGP is introduced to study and understand the differences in the sources of data. The proposed approach is demonstrated on and analyzed through two mathematical (representative parabola problem, 2D Ackley function) and two materials science (design of FeCrAl and SmCoFe alloys) case studies. From the case studies, it is observed that compared to using single-source and source unaware ML models, the proposed multi-source data fusion framework can provide better predictions for sparse-data problems, interpretability regarding the sources, and enhanced modeling capabilities by taking advantage of the correlations and relationships among different sources.
@@
Keywords Latent Variable Gaussian Process, Gaussian Process Regression, Multi-Source Modeling, Data Fusion, Interpretable Artificial Intelligence, Uncertainity Quantification, Probabilistic Machine Learning
@@ 1 Introduction
To tackle the aforementioned challenges, the need for a multi-source data fusion that considers source variabilities becomes evident. By integrating data from diverse sources, multi-source data fusion modeling with interpretable ML can provide a more comprehensive and robust understanding of material behaviors, leading to faster discovery of novel materials. Specifically for experimental data sources, the data-driven ML approach can consider the inherent complexities and uncertainties associated with each experimental data resource through extraction of meaningful patterns and insights from different experimental conditions. This factor effectively mitigates the data availability and quality challenges. To achieve this, a Gaussian Process (GP) based multi-source modeling through Latent Variable Gaussian Process (LVGP) [zhang2020latent] is proposed to fuse numerous information (data) sources together to obtain a single representative interpretable prediction model. LVGP is novel ML model that allows incorporation of qualitative (categorical) variables in addition to quantitative (continuous) variables into surrogate modeling and has been implemented in variety of fields and applications such as adaptive materials design [iyer2020data, wang2020featureless, zhang2020bayesian], non-hierarchical interpretable multi-fidelity modeling [CHEN2024116773], and first-ever mixed-variable (continuous and categorical) metamodel-based global sensitivity analysis [comlek2023mixed]. Therefore, throughout this research, a single multi-source ML model that incorporates engineering system parameters as numerical (and categorical if there are any) variables while incorporating each data source as a qualitative variable is developed and studied. The qualitative information sources are incorporated into the LVGP model by implicitly mapping each qualitative variable into a low-dimensional quantitative latent space through statistical inference techniques. Specifically, as the influence of any qualitative variable on any quantitative response must be always due to some underlying (and possibly high-dimensional) quantitative physical variables, the latent variable approach provides physically meaningful dimension reduction. Here, the term “physically-meaningful” can be associated with explaining the cause-effect relationships between the information sources and the system response. Therefore, the qualitative representation of information sources indirectly accounts for the variability they bring through known and unknown underlying processing variables, which in return provides further interpretability regarding the information sources and their relationships on the system of interest.
The outline of the paper is as follows: In Section 2, descriptions of the ML models that are used for data fusion, GP and LVGP, are given and multi-source data fusion through LVGP is introduced. Next, two exemplary mathematical representative problems in Section 3 are demonstrated to highlight the advantages of using LVGP for multi-source information modeling. In Sections 4 and 5, the proposed multi-source model is applied on two alloy systems (FeCrAl and SmCoFe) to illustrate its benefits for material science research. Finally, the work is concluded through Section 6.
@@ 2.2 Review of Latent Variable Gaussian Process
Due to the nature of the GP correlation functions, it is not plausible to directly include qualitative variables into the modelling since the differences between qualitative variables become unclear without a well-defined distance metric. Recently, Latent Variable Gaussian Process (LVGP) has been developed to integrate qualitative variables into GP modeling through quantitative latent variable representation of these variables to address this difficulty [zhang2020latent].
Consider a new input space, $\mathbf{b}=[\mathbf{x^{T}},\mathbf{t^{T}}]$ where both quantitative, $\mathbf{x}=[x_{1},x_{2},..x_{m}]\epsilon R^{m}$, and qualitative $\mathbf{t}=[t_{1},t_{2},...,t_{q}]\epsilon R^{q}$ input variables are included in this new domain. Each qualitative variable can have $j$ number of options (levels), i.e $t_{i}=[l_{1}(t_{i}),l_{2}(t_{i}),...,l_{j}(t_{i})]$ where $i=1,2,...,q$. For a given response of interest, $y(b)$, the influence of each qualitative variable can be represented by an underlying quantitative (numerical) space, $v_{t_{i}}(l_{j})=[v_{1}(l_{1}),v_{2}(l_{1}),...,v_{k}(l_{j})]\epsilon R^{k}$. This underlying quantitative space can possibly be undiscovered, unknown, complex, or high-dimensional. As a result, the main idea of LVGP lies in learning a low-dimensional quantitative latent space to approximate the original underlying space through statistical inference using the available input-output data. According to the original LVGP paper, a two-dimensional latent variable vector is sufficient to express the influence of the qualitative variables on the response of interest [zhang2020latent]. Therefore, qualitative variable, $t_{i}$ can be represented as $\mathbf{z}(t_{i})=[z_{1}(l_{1}),z_{2}(l_{1}),z_{1}(l_{2}),z_{2}(l_{2}),...,z_{1}(l_{j}),z_{2}(l_{j})]$, where each level $l_{j}(t_{i})$ is represented with a 2D latent vector $[z_{1}(l_{j}),z_{2}(l_{j})]$ Thus, the previously defined mixed-variable input space, $\mathbf{b}=[\mathbf{x}^{T},\mathbf{t}^{T}]$, then becomes $\mathbf{w}=[\mathbf{x}^{T},\mathbf{z(t)}{{}^{T}}]\epsilon R^{m+2\times q}$. With the newly defined input space, the previously specified GP correlation function (Equation 2) is updates as,
@@ 2.3 Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process
As previously mentioned, the latent variable approach provides physics-based dimension reduction that could potentially explain the "cause-effect" relationship between the inputs (the information sources) and outputs (response). Therefore, the proposed model can give us physical insights regarding the information sources by analyzing the spatial locations of the variables (sources) in the latent space. Here, the locations of the latent variables could suggest that the variables that are closer in the latent space can have a similar impact on the response. Moreover, for the scenarios where no prior knowledge exists regarding information sources, the LVGP model can help us extract and reveal the relationships between the sources. To understand the underlying similarities and differences between information sources, a dissimilarity metric ($D$) that quantifies the dissimilarities between the sources through a distance metric is introduced. This dissimilarity metric is measured through the Euclidean distance of the data source $(z^{j})$ from the reference data source $(z^{*})$ in the latent space. The metric is given in Equation 7. For a given LVGP model, the latent variables are constrained to be between $z=[-3,3]$ as this was the typical practice in previous applications of LVGP and the reference data source will be placed at $z^{*}=[0,0]$ in the latent space. Therefore, the most dissimilar information source from the reference source is going to be placed at [±$\pm 3,\pm 3$], making $z^{max}=[\pm 3,\pm 3]$. Based on this, the maximum distance between the reference and given information source can be $||z^{max}-z^{*}||=3\sqrt{2}$. Thus, a normalized unit of dissimilarity ($D$) can be calculated through Equation 8. With this quantified metric, the similarities and differences between information sources can be understood, which could be further used in downstream tasks such as anomaly detection and fault identification of sources.
@@ 3.1 Parabola Problem
Next, Gaussian Process and Latent Variable Gaussian Process (detailed as GP and LVGP in the previous sections) are leveraged to model the response surface of the representative parabola problem ($y$). For the GP model training, only the quantitative (numerical) input variable ($x$) is used to learn the response. In this case, the model is unaware of the sources where the data is generated from. On the other hand, a separate qualitative variable that denotes different information sources is included into the training of the LVGP model, in addition to the quantitative variable ($x$), making the surrogate model aware of the different sources that the data is generated from. The black points marked on Figure 3a displays the quantiative training points used to train the models. A well-known ML model predictive performance metric, Normalized Root Mean Squared Error (NRMSE) is utilized to compare the training and testing performances of these models. Table 1 shows the performance results comparing the two ML models. The low NRMSE value for LVGP clearly shows that incorporating the information source as a separate variable into data fusion ML models yield a better predictive model. Furthermore, the parity plot of the two models (Figure 3b& c), comparing the model predictions and actual values, clearly demonstrates the advantage of incorporating and differentiating the source information into the modeling. Finally, Figure 3d & e shows the predictions made solely on the Ground Source, which once again clearly demonstrates the predictive capabilities of LVGP. In the context of engineering and scientific applications, considering cases where conducting experiments are costly or one has only access to a single experimental resource, obtaining an accurate predictive model for that specific experiment can be highly beneficial for advanced design and development.
@@
The proposed method is next demonstrated on a well-known mathematical function that is used in optimization studies. Ackley’s function is a customiziable multi-dimensional function that has been implemented in numerous optimization applications for testing and validation purposes [ackley2012connectionist, ackley2]. Here, 2D Ackley’s function, Equation 10, is implemented to further demonstrate the proposed interpretable multi-source data fusion method. In addition to the original 2D function, three (3) other variations of Ackley’s function are generated from the original function as different sources of information. Similar to the parabola problem, the original function is named as the Ground Source and the remaining 3 functions are named as Perturbed Sources 1, 2, and 3. The formulations of the functions are found in Supplementary Table 2 and their surface of the response, $z$, is shown in Figure 4a.
@@ 4.1 Overview
Despite their superior oxidation resistance, there is one major hindering aspect of FeCrAl alloys that warrants further study. FeCrAl alloys undergo aging embrittlement which makes them vulnerable to mechanical perturbation and failure. The embrittlement is majorly caused by the Cr-rich $\alpha^{{}^{\prime}}$-phase precipitation in the Fe-rich matrix. The formation of such $\alpha^{{}^{\prime}}$ phases increases the overall hardness of the alloy and makes them brittle.
The thermal age hardening phenomenon is driven by the composition of the FeCrAl alloy, temperature, and duration of the thermal aging process. The hardness change of the alloy at different stages of thermal aging process is measured and used as an indication of the embrittlement. With the nuance of thermal aging phenomenon in FeCrAl alloys, material designers have the task of optimizing the alloy composition for maximum oxidation resistance and minimum embrittlement. The levers that material designers have at their disposal is the composition of the FeCrAl alloys including dopants that can suppress the formation of $\alpha^{{}^{\prime}}$ phases.
@@ 4.4 Results and Discussion
Within this study, (1) the generalizability of the models, and (2) the performances of the models with respect to the experimental in-house information source that can be controlled is investigated. The main goal of (1) is to observe the usefulness of the models when one has control to multiple sources of information to perform further experimentations, whereas the main goal of (2) is to observe which of the models is more reliable when one has further control over a source of information. Out of all 380 available samples, 344 samples are used to observe the generalizability of models through cross validation (CV) and set aside 36 GE Data points as a separate testing set to demonstrate the prediction capabilities of the models with respect to the information source that can be control and perform further experiments on. For the remaining 344 samples, 5-fold cross validation (around 275 training and 69 validation samples per fold) is employed to evaluate the generalizability of both models. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. Then, both models are trained using the 344 samples and tested on the separate 36 GE testing data. The training, validation and testing samples were held identical for both models. The results of this study are shown in Section 4.4.1. In addition to comparing two ML models on the multi-source data, LVGP is further compared with a separate GP model trained only on GE Data. For ease of notation, this GP model is denoted as GP-GE. GP-GE was only trained on GE Data samples available in the remaining 344 sample set (142 samples). Both LVGP and GP-GE are tested on the same separate 36 GE Data samples. The results of this study are shown in Section 4.4.2.
@@ 4.4.1 Multi-Source Data Fusion Comparison: LVGP and GP
To strengthen our case of using multi-source modeling through LVGP, response surface plots from both the GP and LVGP models (with GE Data as reference) are also provided through Panels A and B of Figure 6. The second and third row depict the mean and standard deviation of the hardness changes across variation of the Cr (12%-24%) and Al (0%-7%) and the fourth row depicts the combined mean and standard deviation in a 3D representation. The model validity is assessed by focusing on the mean and variation behavior of the response surface. It is noted that the response surface of the hardness changes are highly undulated with sharp peaks and valleys with the GP model. On the other hand, using the GE source as ground truth for prediction, the response surface is much smoother and more well-behaved with the LVGP model. One can argue that the unsmooth behavior is caused by the varying level of noises coming from different sources in the data set when fitting all of them to one single GP model. On the other hand, LVGP has the capability of smoothing out the curves from multiple sources after the introduction of the latent variables. From a material science perspective, it is very unlikely that phenomenon driving the hardness changes across the composition are highly transient and acute. This important observation highlights the need of accounting for known and unknown underlying physical parameters when data from numerous sources are combined for data fusion. The samples collected from different papers might have had differences in manufacturing process, thermal aging procedure, and testing technique that are not accounted for in a single source model. Furthermore, the downstream tasks of optimization, sensitivity analysis, uncertainty quantification would also be severely affected by such distinctions and sharp undulations of the surrogate models, thus making the case of employing multi-source modeling through LVGP once again. Finally, LVGP has the capability of providing source-specific predictions, which is not plausible with a single GP model that cannot differentiate between the sources. The response surface predictions provided by LVGP model for the remaining sources of information are provided in Supplementary Figure 7. This property provided by the LVGP can be highly advantageous for the cases when one has control over multiple sources of information.

### v4

@@ Abstract
With the advent of artificial intelligence and machine learning, various domains of science and engineering communities have leveraged data-driven surrogates to model complex systems through fusing numerous sources of information (data) from published papers, patents, open repositories, or other resources. However, not much attention has been paid to the differences in quality and comprehensiveness of the known and unknown underlying physical parameters of the information sources, which could have downstream implications during system optimization. Additionally, existing methods cannot fuse multi-source data into a single predictive model. Towards resolving this issue, a multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed. The individual data sources are tagged as a characteristic categorical variable that are mapped into a physically interpretable latent space, allowing the development of source-aware data fusion modeling. Additionally, a dissimilarity metric based on the latent variables of LVGP is introduced to study and understand the differences in the sources of data. The proposed approach is demonstrated on and analyzed through two mathematical and two materials science case studies. From the case studies, it is observed that compared to using single-source and source unaware machine learning models, the proposed multi-source data fusion framework can provide better predictions for sparse-data problems.
@@
Keywords Latent Variable Gaussian Process, Gaussian Process Regression, Multi-Source Modeling, Data Fusion, Interpretable Artificial Intelligence, Uncertainty Quantification, Probabilistic Machine Learning
@@ 1 Introduction
To tackle the aforementioned challenges, the need for a multi-source data fusion that considers source variabilities becomes evident. In this work, a Gaussian Process (GP) based multi-source data fusion modeling through Latent Variable Gaussian Process (LVGP) [zhang2020latent] is proposed to fuse numerous information (data) sources together to obtain a single representative interpretable prediction model. LVGP is a novel ML model that allows incorporation of qualitative (categorical) variables in addition to quantitative (continuous) variables into surrogate modeling. Throughout this research, in addition to incorporating engineering system parameters as numerical (and categorical if there are any), LVGP is used to incorporate each data source as a categorical variable into the data fusion modeling. The categorical source variables are incorporated into the data fusion model by implicitly mapping each variable onto a physically-meaningful low-dimensional numerical latent space through statistical inference techniques. Here, the term “physically-meaningful” can be associated with explaining the cause-effect relationships between the information sources and the system response. Therefore, the categorical representation of the information sources can indirectly account for the variability that the sources bring through known and unknown underlying processing variables, which in return provides interpretability regarding the information sources and their relationships on the system of interest. Furthermore, through extraction of meaningful spatial patterns from the latent space, data availability and quality challenges can be mitigating by carefully selecting which sources to be used for modeling. As a result, the innovation contributions of this work are as follows: (1) a single source-aware multi-source data fusion model, (2) interpretable latent space that explains the relationships between sources, (3) a dissimilarity metric on the latent space that can be used for targeted source selection for improved data fusion modeling, and (4) application of the proposed method on real engineering problems.
The outline of the paper is as follows: In Section 2, descriptions of the ML models that are used for data fusion, GP and LVGP, are given and multi-source data fusion through LVGP is introduced. Section 3 demonstrates two exemplary mathematical representative problems to highlight the advantages of using LVGP for multi-source information modeling. In Sections 4 and 5, the proposed multi-source model is applied on two alloy systems (FeCrAl and SmCoFe) to illustrate its benefits for material science research. Finally, the work is concluded through Section 6.
@@ 2.2 Review of Latent Variable Gaussian Process
Due to the nature of the GP correlation functions, it is not plausible to directly include qualitative variables into the modelling since the differences between qualitative variables become unclear without a well-defined distance metric. Recently, Latent Variable Gaussian Process (LVGP) has been developed to integrate qualitative variables into GP modeling through quantitative latent variable representation of these variables to address this difficulty [zhang2020latent]. Many prior works have demonstrated the remarkable modeling capabilities of LVGP for different engineering applications and techniques with mixed-variable (continuous and categorical) domains including but not limited to adaptive materials design [iyer2020data, wang2020featureless, zhang2020bayesian, comlek2023rapid, wang2021data], non-hierarchical interpretable multi-fidelity modeling [CHEN2024116773], and mixed-variable (continuous and categorical) global sensitivity analysis [comlek2023mixed].
To briefly review the method, consider a new input space, $\mathbf{b}=[\mathbf{x^{T}},\mathbf{t^{T}}]$ where both quantitative, $\mathbf{x}=[x_{1},x_{2},..x_{m}]\epsilon R^{m}$, and qualitative $\mathbf{t}=[t_{1},t_{2},...,t_{q}]\epsilon R^{q}$ input variables are included in this new domain. Each qualitative variable can have $j$ number of options (levels), i.e $t_{i}=[l_{1}(t_{i}),l_{2}(t_{i}),...,l_{j}(t_{i})]$ where $i=1,2,...,q$. For a given response of interest, $y(b)$, the influence of each qualitative variable can be represented by an underlying quantitative (numerical) space, $v_{t_{i}}(l_{j})=[v_{1}(l_{1}),v_{2}(l_{1}),...,v_{k}(l_{j})]\epsilon R^{k}$. This underlying quantitative space can possibly be undiscovered, unknown, complex, or high-dimensional. As a result, the main idea of LVGP lies in learning a low-dimensional quantitative latent space to approximate the original underlying space through statistical inference using the available input-output data. Specifically, it is assumed that this low dimensional space can capture the joint effects of these underlying qualitative variables on the response of interest according to sufficient dimension reduction arguments [li1991sliced, cook2005sufficient]. As such, according to the authors of the original LVGP paper [zhang2020latent], a two-dimensional latent variable vector is sufficient to express the influence of the qualitative variables on the response. Therefore, qualitative variable, $t_{i}$ can be represented as $\mathbf{z}(t_{i})=[z_{1}(l_{1}),z_{2}(l_{1}),z_{1}(l_{2}),z_{2}(l_{2}),...,z_{1}(l_{j}),z_{2}(l_{j})]$, where each level $l_{j}(t_{i})$ is represented with a 2D latent vector $[z_{1}(l_{j}),z_{2}(l_{j})]$. Thus, the previously defined mixed-variable input space, $\mathbf{b}=[\mathbf{x}^{T},\mathbf{t}^{T}]$, then becomes $\mathbf{w}=[\mathbf{x}^{T},\mathbf{z(t)}{{}^{T}}]\epsilon R^{m+2\times q}$. With the newly defined input space, the GP correlation function (Equation 2) is updated as,
@@ 2.3 Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process
As previously mentioned, the latent variable approach provides physics-based dimension reduction that could potentially explain the "cause-effect" relationship between the inputs (the information sources) and outputs (response). Therefore, the proposed model can give us physical insights regarding the information sources by analyzing the spatial locations of the variables (sources) in the latent space. Here, the locations of the latent variables could suggest that the variables that are closer in the latent space can have a similar impact on the response. Moreover, for the scenarios where no prior knowledge exists regarding information sources, the LVGP model can help us extract and reveal the relationships between the sources. To understand the underlying similarities and differences between information sources, a dissimilarity metric ($D$) that quantifies the dissimilarities between the sources through a distance metric is introduced. This dissimilarity metric is measured through the Euclidean distance of the data source $(z^{j})$ from the reference data source $(z^{*})$ in the latent space. The metric is given in Equation 7. For a given LVGP model, the latent variables are constrained to be between $z=[-3,3]$ as this was the typical practice in previous applications of LVGP and the reference data source will be placed at $z^{*}=[0,0]$ in the latent space. Therefore, the most dissimilar information source from the reference source is going to be placed at [±$\pm 3,\pm 3$], making $z^{max}=[\pm 3,\pm 3]$. Based on this, the maximum distance between the reference and given information source can be $||z^{max}-z^{*}||=3\sqrt{2}$. Thus, a normalized unit of dissimilarity ($D$) can be calculated through Equation 8. With this quantified metric, the similarities and differences between information sources can be understood, which could be further used in downstream tasks such as anomaly detection and fault identification of sources. Finally, although the choice of reference source is determined before the LVGP model training, the reference source choice does not influence the spatial relationships obtained from the latent space. As a result, independent of reference source selection, the same spatial relationship between the sources, hence same $D$ values between sources will be obtained for any reference source selection. Thus, the reference source selection is user, goal, and application specific.
@@ 3.1 Parabola Problem
Next, Gaussian Process and Latent Variable Gaussian Process (detailed as GP and LVGP in the previous sections) are leveraged to model the response surface of the representative parabola problem ($y$). For the GP model training, only the quantitative (numerical) input variable ($x$) is used to learn the response. In this case, the model is unaware of the sources where the data is generated from. On the other hand, a separate qualitative variable that denotes different information sources is included into the training of the LVGP model, in addition to the quantitative variable ($x$), making the surrogate model aware of the different sources that the data is generated from. The black points marked on Figure 3a displays the quantitative training points used to train the models. A well-known ML model predictive performance metric, Normalized Root Mean Squared Error (NRMSE) is utilized to compare the training and testing performances of these models. Table 1 shows the performance results comparing the two ML models. The low NRMSE value for LVGP clearly shows that incorporating the information source as a separate variable into data fusion ML models yield a better predictive model. Furthermore, the parity plot of the two models (Figure 3b& c), comparing the model predictions and actual values, clearly demonstrates the advantage of incorporating and differentiating the source information into the modeling. Finally, Figure 3d & e shows the predictions made solely on the Ground Source, which once again clearly demonstrates the predictive capabilities of LVGP. In the context of engineering and scientific applications, considering cases where conducting experiments are costly or one has only access to a single experimental resource, obtaining an accurate predictive model for that specific experiment can be highly beneficial for advanced design and development.
@@
The proposed method is next demonstrated on a well-known mathematical function that is used in optimization studies. Ackley’s function is a customizable multi-dimensional function that has been implemented in numerous optimization applications for testing and validation purposes [ackley2012connectionist, ackley2]. Here, 2D Ackley’s function, Equation 10, is implemented to further demonstrate the proposed interpretable multi-source data fusion method. In addition to the original 2D function, three (3) other variations of Ackley’s function are generated from the original function as different sources of information. Similar to the parabola problem, the original function is named as the Ground Source and the remaining 3 functions are named as Perturbed Sources 1, 2, and 3. The formulations of the functions are found in Supplementary Table 2 and their surface of the response, $z$, is shown in Figure 4a.
@@ 4.1 Overview
Despite their superior oxidation resistance measured through change in specific mass of the alloy coupon, there is one major hindering aspect of FeCrAl alloys that warrants further study. FeCrAl alloys undergo aging embrittlement which makes them vulnerable to mechanical perturbation and failure. The embrittlement is majorly caused by the Cr-rich $\alpha^{{}^{\prime}}$-phase precipitation in the Fe-rich matrix. The formation of such $\alpha^{{}^{\prime}}$ phases increases the overall hardness of the alloy and makes them brittle.
The thermal age hardening phenomenon is driven by the composition of the FeCrAl alloy, temperature, and duration of the thermal aging process. The hardness change measured through Vicker’s hardness of the alloy at different stages of thermal aging process is studied and used as an indication of the embrittlement. With the nuance of thermal aging phenomenon in FeCrAl alloys, material designers have the task of optimizing the alloy composition for maximum oxidation resistance and minimum embrittlement. The levers that material designers have at their disposal is the composition of the FeCrAl alloys including dopants that can suppress the formation of $\alpha^{{}^{\prime}}$ phases.
@@ 4.4 Results and Discussion
Within this study, (1) the generalizability of the models, and (2) the performances of the models with respect to the experimental in-house information source that can be controlled is investigated. The main goal of (1) is to observe the usefulness of the models when one has control to multiple sources of information to perform further experimentations, whereas the main goal of (2) is to observe which of the models is more reliable when one has further control over a source of information. Out of all 380 available samples, 344 samples are used to observe the generalizability of models through cross validation (CV) and set aside 36 GE Data points as a separate testing set to demonstrate the prediction capabilities of the models with respect to the information source that can be control and perform further experiments on. For the remaining 344 samples, 5-fold cross validation (around 275 training and 69 validation samples per fold) is employed to evaluate the generalizability of both models. For CV tasks in LVGP, stratified sampling is used to make sure datasets used for training and testing represent the overall distribution of the information sources. Then, both models are trained using the 344 samples and tested on the separate 36 GE testing data. The training, validation and testing samples were held identical for both models. The results of this study are shown in Section 4.4.1. In addition to comparing two ML models on the multi-source data, LVGP is further compared with a separate GP model trained only on GE Data. For ease of notation, this GP model is denoted as GP-GE. GP-GE was only trained on GE Data samples available in the remaining 344 sample set (142 samples), while the previously trained LVGP model on the 344 sample set is used for comparison. Both LVGP and GP-GE are tested on the same separate 36 GE Data samples. The results of this study are shown in Section 4.4.2.
@@ 4.4.1 Multi-Source Data Fusion Comparison: LVGP and GP
To strengthen our case of using multi-source modeling through LVGP, response surface plots from both the GP and LVGP models (with GE Data as reference) are also provided through Panels A and B of Figure 6. The second and third row depict the mean and standard deviation of the hardness changes across variation of the Cr (12%-24%) and Al (0%-7%) and the fourth row depicts the combined mean and standard deviation in a 3D representation. The model validity is assessed by focusing on the mean and variation behavior of the response surface. It is noted that the response surface of the hardness changes are highly undulated with sharp peaks and valleys with the GP model. On the other hand, using the GE source as ground truth for prediction, the response surface is much smoother and more well-behaved with the LVGP model. One can argue that the unsmooth behavior is caused by the varying level of noises coming from different sources in the dataset when fitting all of them to one single GP model. On the other hand, LVGP has the capability of smoothing out the curves from multiple sources after the introduction of the latent variables. From a material science perspective, it is very unlikely that phenomenon driving the hardness changes across the composition are highly transient and acute. This important observation highlights the need of accounting for known and unknown underlying physical parameters when data from numerous sources are combined for data fusion. The samples collected from different papers might have had differences in manufacturing process, thermal aging procedure, and testing technique that are not accounted for in a single source model. Furthermore, the downstream tasks of optimization, sensitivity analysis, uncertainty quantification would also be severely affected by such distinctions and sharp undulations of the surrogate models, thus making the case of employing multi-source modeling through LVGP once again. Finally, LVGP has the capability of providing source-specific predictions, which is not plausible with a single GP model that cannot differentiate between the sources. The response surface predictions provided by LVGP model for the remaining sources of information are provided in Figure 7. This property provided by the LVGP can be highly advantageous for the cases when one has control over multiple sources of information.